OpenStudy (anonymous):

derivatieve of x^2 secy assuming that the equation determines a differentiable function f suht that y= f(x)

OpenStudy (anonymous):

technically cant be done

OpenStudy (anonymous):

it has to equal something

OpenStudy (anonymous):

x^2 secy must equal something

OpenStudy (anonymous):

= y^2 +1

OpenStudy (anonymous):

x^2 (cosy)^-1 = y^2 +1

OpenStudy (anonymous):

rewritting sec as cos^-1

OpenStudy (anonymous):

oh i used the product rule

OpenStudy (anonymous):

yeh, it is correct

OpenStudy (anonymous):

(2x)(secy) but i dont know whts goes in the next bracket

OpenStudy (anonymous):

because y depends on x , x^2 (cosy)^-1 needs to be differenitated as product

OpenStudy (anonymous):

secytany(x^2)y'

OpenStudy (anonymous):

so then it'll be 2yy' = 2xsecy+x^2secytany*y'

OpenStudy (anonymous):

so LHS: let u= x^2 du/dx = 2x v= (cos(y))^-1 dv/dx = -1 ( cos(y) )^-2 * -sin(y) (dy/dx)

OpenStudy (anonymous):

so the LHS = 2x (cosy)^-1 +(x^2sin(y) (sec(y))^2 ) dy/dx

OpenStudy (anonymous):

and the RHS will equal 2y dy/dx

OpenStudy (anonymous):

\[2xsec(y) +x^2\sin(y)\sec^2(y)\frac{dy}{dx} = 2y \frac{dy}{dx}\]

OpenStudy (anonymous):

solve for the dy/dx

OpenStudy (anonymous):

ok i understand thank you so much